AB and CD are two lines of length 10 cm. and 5 cm. respectively.

We choose point P on line AB and point Q on line CD at random.

Length of CQ=X and length of AP=Y are two independent random variable

a)   P(Y > X)=?

b)  P( X=2. Y=2)=?



1
Expert's answer
2021-01-04T18:55:12-0500

PX(x)=1/5 for 0"\le X \le 5"

PY(y)=1/10 for 0"\le Y \le 10"

P(Y > X)="\int_{x=0}^{5} \int_{y=x}^{10} P_X(x)*P_Y(y) dydx= \int_{x=0}^{5} \int_{y=x}^{10} 1/50 dydx="

"\int_{x=0}^{5} (10-x)/50*dx=-1/50*\int_0^5x*dx+1/5\int_0^51*dx=-5^2/100+5/5=3/4"

P( X=2 and Y=2) = P(X=2)*P(Y=2) = "(\int_2^{2} 1/5*dx)*(\int_2^{2} 1/10*dy)=0"